# Combine and conquer: model averaging for out-of-distribution forecasting

### v1

@@ Abstract
Travel behaviour modellers have an increasingly diverse set of models at their disposal, ranging from traditional econometric structures to models from mathematical psychology and data-driven approaches from machine learning. A key question arises as to how well these different models perform in prediction, especially when considering trips of different characteristics from those used in estimation, i.e. out-of-distribution prediction, and whether better predictions can be obtained by combining insights from the different models. Across two case studies, we show that while data-driven approaches excel in predicting mode choice for trips within the distance bands used in estimation, beyond that range, the picture is fuzzy. To leverage the relative advantages of the different model families and capitalise on the notion that multiple ‘weak’ models can result in more robust models, we put forward the use of a model averaging approach that allocates weights to different model families as a function of the distance between the characteristics of the trip for which predictions are made, and those used in model estimation. Overall, we see that the model averaging approach gives larger weight to models with stronger behavioural or econometric underpinnings the more we move outside the interval of trip distances covered in estimation. Across both case studies, we show that our model averaging approach obtains improved performance both on the estimation and validation data, and crucially also when predicting mode choices for trips of distances outside the range used in estimation.
@@ 1 Introduction
Travel behaviour analysts are increasingly interested in a more diverse set of models, covering econometric structures, models from mathematical psychology, and machine learning techniques. With the typical model evaluation approach focussing on the ability of models to replicate the behaviour in the empirical data as well as possible, it should then come as no surprise that researchers in the travel behaviour field are increasingly attracted by the comparatively good prediction performance of machine learning approaches (cf. Hagenauer and Helbich, 2017).
@@
The aim of this paper is twofold. First, we aim to shed light on the prediction performance of forecasting models with varying levels of behavioural underpinning as a function of the “distance” between the training data and forecasting scenarios for models. Second, we aim to develop an approach that improves prediction performance in forecasts through a model averaging approach that assigns different weights to different models depending on this “distance". In our work, we use trip distance as the key variable, but the approach works with any variable that goes out of distribution, not only physical distance.
@@ 2.1 Generic model averaging approach
In model averaging (see e.g. Hancock et al., 2020), we first estimate $M$ different models on the data, with model $m$ giving a likelihood $L_{n,t,m}\left(\Omega_{m}\right)$ for the choice observed by decision maker $n$ in choice situation $t$, using a vector of estimated model parameters $\Omega_{m}$. Let us further define $L_{n,m}\left(\Omega_{m}\right)$ to be the likelihood for the entire sequence of $T_{n}$ choices observed for person $n$ according to model $m$.
@@
In applying the model to predict choices out-of-sample, an analyst then again requires the predicted choice probabilities from the individual models first. Let us assume that $P_{i,n,f,m}\left(\hat{\Omega}_{m}\right)$ is the probability of person $n$ choosing alternative $i$ in some forecast scenario $f$, according to model $m$, again calculated at the MLEs. Let us further assume that we have worked with model averaging weights at the observation level, i.e. maximising Equation 2 to obtain MLEs for the model weights at the observation level. We can then use these parameters to calculate the model averaging weights for the forecast scenario, i.e. $\pi_{n,f,m}\left(\hat{\Theta}\right)$. The model averaging probability for choosing alternative $i$ in this new setting would then be given by:
@@ 2.2 Model averaging for out-of-distribution prediction
Let us focus on journey distance in particular. Training a model averaging structure where journey distance, say $d_{n,t}$ for person $n$ and trip $t$, is used as a key characteristic in Equation 4 would allow the model to learn how different sub-models are better suited for predicting trips of specific distances within the bounds of the estimation data. However, if the individual sub-models have been estimated on the entire data, then model averaging by definition cannot gain any insights about how the performance might differ across models out-of-distribution, and thus what weights should be given to them outside the range of distances used in estimation. As a result, the weights will simply be an extrapolation based on the estimates for $\gamma_{n,t,m}$ obtained on the training data. For example, imagine that our model would find that the weight assigned to model $m$ increases with distance. If making predictions out-of-distribution, the model would follow the same trend, and for trips with distances below the lower bound of the estimation data, the weight assigned to model $m$ would decrease, while for trips with distances above the upper bound of the estimation data, the weight assigned to model $m$ would increase.
@@ 3.1.1 DECISIONS data
After extensive data cleaning and data enrichment (Tsoleridis et al., 2022), 12,524 trips made by 540 individuals remained. For each trip, individuals travelled by one of six modes: car (48%), bus (15%), rail (5%), taxi (3%), cycling (3%) or walking (26%). The data covers a very wide range of trip distances, going from just under 100 metres to just under $106$ kilometres, with a mean distance of $8.2$ kilometres and a median distance of $4.5$ kilometres.
@@ 3.1.2 LPMC data
3.2 Estimation and validation data
@@
In addition to these divisions, we further retain $20\%$ of any trips, independent of their distance, for out-of-sample validation [^1] The splitting of the sample into estimation, validation (and additionally test data for MLP and XGB) was done at the observation rather than person/household level. This might be seen as a limitation, however, none of the models we used explicitly accommodated any individual-specific effects, and information on IDs was not provided to the MLP and XGB training. We thus believe the impact of the way we have split the data to be negligible.. The resulting division of the data is explained in Figure 1, with the resulting sample sizes shown Table 1, where the larger sample validation size for the two extreme segments is a result of this data being used for validation only. In practical work, analysts would be unlikely to not include all deciles (intervals of 10% width) in the MA estimation work and would likely also use a smaller validation sample - we made these decisions with a view to validating the model idea.
Figure 1: Estimation and validation data
Table 1: Estimation and validation data sample sizes for case studies
@@
Distance (km) Estimation Validation Distance (km) Estimation Validation
@@
out-of-distribution for MA (below, i.e. $d_{min}tod_{a,MA}$) <1.33 0 0 1,253 <0.75 0 0 8,092
out-of-distribution for sub-models (below, i.e. $d_{a,MA}$ to $d_{a,sub}$) 1.33 to 1.95 0 1,026 228 0.75 to 1.11 0 6,508 1,580
Within distribution for MA and sub-models 1.95 to 11.95 5,952 1,560 1.11 to 7.41 38,855 9,696
out-of-distribution for sub-models (above, i.e. $d_{b,sub}$ to $d_{b,MA}$) 11.95 to 19.52 0 993 259 7.41 to 11.22 0 6,481 1,623
out-of-distribution for MA (above, i.e. $d_{b,MA}$ to $d_{max}$) 19.52 to 105.9 0 0 1,253 11.22 to 40.1 0 0 8,108
Total 5,952 2,019 4,553 38,855 12,989 29,099
@@ 3.3 Model types
The following individual models are used in our analysis, combining structures from choice modelling, mathematical psychology and machine learning. Many other model types could have been considered [^2] A notable omission in our list are mixture models such as mixed logit or latent class - any advantages that such models have in estimation largely/completely disappear in prediction as the random heterogeneity needs to be averaged out. The only exception comes with the use of posterior distributions which are not applicable in the context of out-of-sample prediction. - we limited ourselves to this set with a view to demonstrating the potential benefit of our proposed approach.
@@
Our NL models (cf. Daly and Zachary, 1978) use a full set (i.e. $J-1$) of alternative specific constants (ASC). For continuous attributes, we use a linear plus logarithmic specification, i.e. $V=...+\beta_{x}x+\beta_{log-x}log(x)$ to capture non-linearity in sensitivities. For the DECISIONS data, we use three nests, grouping car together with taxi, bus with train, and walking with cycling. For the LPMC data, we use two nests, grouping car together with bus, and walking together with cycling, but this structure ended up collapsing to a Multinomial Logit (MNL) model.
@@
This model comprises an input layer with input nodes, one or more hidden layers with hidden nodes, and an output layer with output nodes. In this model, signals propagate forward through the links connecting the nodes. The links have numeric weights $w$, which are learned from the data. At each link, the weights are multiplied by the input value from the previous nodes. At the node, the sum of the inputs is taken, and an activation function is applied. The outcome is propagated to the nodes in the next layer. We use tanh activation functions in all hidden layers, except the final one where a linear layer is used. In the output layer, a Softmax function (i.e. a logit) is applied to produce choice probabilities for each alternative.
@@
The XGB model comprises a series of sequentially applied decision trees. A decision tree is a sequence of simple IF-THEN rules, optimised to classify data accurately. In the XGB model, each decision tree in the series ‘corrects’ the mispredictions of the models before it. This process is referred to as ‘boosting’. The term ‘gradient’ refers to the fact that each successive tree is fitted using the gradient of the loss function—effectively via gradient descent. Unlike standard Gradient Boosting Machines (GBM), XGBoost includes further optimisations, such as second-order gradient information and advanced regularisation, to improve speed and mitigate overfitting.
@@
The NL and DFT models were coded and estimated using Apollo (cf. Hess and Palma, 2019), while the MLP and XGB models were implemented in Python using the SciKit-learn (Pedregosa et al., 2011) and XGBoost (Chen and Guestrin, 2016) libraries respectively. A hyperparameter search was conducted for both models to optimise their performance.
@@
For the MLP, the hyperparameter search considered the number of hidden layers, the number of nodes per layer, the learning rate, and the L2 regularisation strength. The number of hidden layers and nodes per layer influences the model’s capacity to learn complex patterns, while the learning rate controls how quickly the model updates its weights during training. L2 regularisation helps prevent overfitting by penalising large weight values. Since the best-performing hyperparameters were highly similar across both datasets, we opted to use a single configuration. The best-performing configuration found was two hidden layers with 30 nodes each, a learning rate of 0.001, and an L2 regularisation strength of 0.1. Because each training run of an MLP can yield slightly different outcomes, we trained the model 100 times using the best-performing configuration and averaged their predictions to account for this variability.
For the XGBoost model, the hyperparameter search focused on selecting a subset of features for each tree split, determining the minimum loss reduction needed for further node splitting, setting the maximum tree depth, and specifying the minimum sum of instance weights required in a child node. The first controls the diversity of the model by randomly selecting a fraction of features at each split. The second acts as a pruning mechanism by restricting splits that do not provide sufficient improvement. The third helps balance model complexity and the risk of overfitting. The fourth serves as a regularisation measure, preventing excessively complex splits by enforcing a minimum requirement for node creation. The final parameter determines the fraction of training samples used in each boosting iteration. Similar to the MLPs, since the optimal hyperparameters were nearly identical for both datasets, we opted to use a single configuration. The best-performing values selected were a feature sampling rate of 0.7, a pruning threshold of 2, a maximum tree depth of 2, a minimum child node weight requirement of 1, and a subsampling rate of 1333This is rather shallow, especially in comparison with Hillel (2021), but was the outcome of the hyperperformance tunning on the subset of the data we used.. As with the MLP model, we repeated the training 100 times using the optimal configuration and averaged their predictions to account for variability in the model training. This is common practice, as when training MLPs, a substantial share gets stuck in inferior solutions.
@@ 3.4 Model averaging: specification and estimation
Unlike the MLP models implemented for the training on the data sets (see Section 3.3), the MLP for the model averaging is implemented in PyTorch Paszke et al. (2019). This is needed because the loss function (see equation (2)) requires a custom implementation, which is not possible in SciKit-learn. To find the best hyperparameters, hyperparameter tuning was conducted using the following parameters: batch size, number of hidden layers and nodes, learning rate and L2 regularisation. For the DECISIONS data set, the best-performing model has two hidden layers with 60 nodes each, a batch size of 128, a learning rate of 0.005, and an L2 regularisation strength of 0.0001. For the LPMC data set, the best-performing model has two hidden layers with 60 nodes each, a batch size of 64, a learning rate of 0.001, and an L2 regularisation strength of 0.0001.
We repeated the training 100 times using the optimal configuration. As we aim for high prediction performance, we retained only the top 20 percent of models [^4] This selection was based on the log-likelihood. For this, we used the full data, i.e. training and test data, but excluded the validation data., discarding those that became trapped in poor local maxima. This approach, thus, differs slightly from the one used for the training of the data-driven models, where we also repeated training 100 times but averaged predictions across all models to encourage diversity between models.
Figure 2 shows the structure for our model averaging approach. As shown in the figure, the model averaging MLP only takes the trip distance as input to predict the weights given to each sub-model. The model-averaging probability is computed as a weighted sum, combining the predictions for the trip by each sub-model and the weights assigned to each sub-model by the model averaging MLP.

### v2

@@ Abstract
Travel behaviour modellers have an increasingly diverse set of models at their disposal, ranging from traditional econometric structures to models from mathematical psychology and data-driven approaches from machine learning. A key question arises as to how well these different models perform in prediction, especially when considering trips of different characteristics from those used in estimation, i.e. out-of-distribution prediction, and whether better predictions can be obtained by combining insights from the different models. Across two case studies, we show that while data-driven approaches excel in predicting mode choice for trips within the distance bands used in estimation, beyond that range, the picture is fuzzy. To leverage the relative advantages of the different model families and capitalise on the notion that multiple ‘weak’ models can result in more robust models, we put forward the use of a model averaging approach that allocates weights to different model families as a function of the distance between the characteristics of the trip for which predictions are made, and those used in model estimation - in our case, we use trip distance as the sole characteristic for this. Overall, we see that the model averaging approach gives larger weight to models with stronger behavioural or econometric underpinnings the more we move outside the interval of trip distances covered in estimation. Across both case studies, we show that our model averaging approach obtains improved performance both on the estimation and test data, and crucially also when predicting mode choices for trips of distances outside the range used in estimation.
@@ 1 Introduction
Travel behaviour analysts are increasingly interested in a more diverse set of models, covering not just econometric structures, but also models from mathematical psychology (e.g. Hancock et al., 2021, 2020a), and increasingly machine learning techniques (e.g. Alwosheel et al., 2021; Wong and Farooq, 2021; Sfeir et al., 2022; Yan et al., 2025; Wang et al., 2020). This is especially true given the growing reliance on highly granualar data on travel behaviour, for example that collected using GPS tracking (e.g. Liu et al., 2024; Tsoleridis et al., 2025).
With the typical model evaluation approach focussing on the ability of models to replicate the behaviour in the empirical data as well as possible, it should then come as no surprise that researchers in the travel behaviour field are increasingly attracted by the comparatively good prediction performance of machine learning approaches (cf. Hagenauer and Helbich, 2017).
@@
The aim of this paper is twofold. First, we aim to shed light on the prediction performance of forecasting models with varying levels of behavioural underpinning as a function of the “distance” between the training data and forecasting scenarios for models. Second, we aim to develop an approach that improves prediction performance in forecasts through a model averaging approach that assigns different weights to different models depending on this “distance". In our work, we use trip distance as the key variable, but the approach works with any variable that goes out of distribution, not only geographic distance.
@@ 2.1 Generic model averaging approach
In model averaging (see e.g. Hancock et al., 2020b), which can be viewed as a form of stacking, first introduced in the seminal work of Wolpert (Wolpert, 1992), which established the conceptual basis for ensemble methods combining multiple predictive models. Specifically, we first estimate $M$ different models on the data, with model $m$ giving a likelihood $L_{n,t,m}\left(\Omega_{m}\right)$ for the choice observed by decision maker $n$ in choice situation $t$, using a vector of estimated model parameters $\Omega_{m}$. Let us further define $L_{n,m}\left(\Omega_{m}\right)$ to be the likelihood for the entire sequence of $T_{n}$ choices observed for person $n$ according to model $m$.
@@
[…]
@@ 2.2 Model averaging for out-of-distribution prediction
In the present paper, we focus on trip distance in particular. This is to some extent a pragmatic choice given the proof-of-concept nature of our work. It allows us to very easily split the data and define what constitutes within-distribution and out-of-distribution, with a view to demonstrating the performance of the model averaging approach.
In addition however, trip distance has a number of appealing characteristics in this context. First, distance is not typically used as an explanatory variable in the mode choice models themselves, which instead rely on the closely correlated variables such as time and cost. Second, distance can explain much of the similarity between trips, along multiple dimensions at the same time. This way, the model averaging structure can learn which model works well for what trip distance, without having to rely on multiple covariates, given that this single variable condenses other information into a single factor.
Training a model averaging structure where journey distance, say $d_{n,t}$ for person $n$ and trip $t$, is used as a key characteristic in Equation 4 would allow the model to learn how different sub-models are better suited for predicting trips of specific distances within the bounds of the estimation data. However, if the individual sub-models have been estimated on the entire data, then model averaging by definition cannot gain any insights about how the performance might differ across models out-of-distribution, and thus what weights should be given to them outside the range of distances used in estimation. As a result, the weights will simply be an extrapolation based on the estimates for $\gamma_{n,t,m}$ obtained on the training data. For example, imagine that our model would find that the weight assigned to model $m$ increases with distance. If making predictions out-of-distribution, the model would follow the same trend, and for trips with distances below the lower bound of the estimation data, the weight assigned to model $m$ would decrease, while for trips with distances above the upper bound of the estimation data, the weight assigned to model $m$ would increase.
@@ 3.1.1 DECISIONS data
After extensive data cleaning and data enrichment (Tsoleridis et al., 2022), 10,990 trips made by 415 individuals remained. For each trip, individuals travelled by one of six modes: car (48%), bus (14%), rail (5%), taxi (3%), cycling (4%) or walking (26%). The data covers a very wide range of trip distances, going from just under 100 metres to just under $106$ kilometres, with a mean distance of $8.2$ kilometres and a median distance of $4.5$ kilometres.
@@ 3.1.2 LPMC data
3.2 Estimation and test data
@@
In addition to these divisions, we further retain $20\%$ of any trips, independent of their distance, for out-of-sample test. The splitting of the sample into estimation, test (and additionally validation data for MLP and XGB) was done at the person/household level, so as to avoid data leakage. In other words, each person only contributed data to either the estimate data or the test data. For those individuals who contributed to the estimation data, their observations in distance segments 1 and 10 were thus not used in our analysis. The resulting division of the data is explained in Figure 1, with the resulting sample sizes shown Table 1. In practical work, analysts would be unlikely to not include all deciles (intervals of 10% width) in the MA estimation work - we made these decisions with a view to validating the model idea. Similarly, a real-world application would not set aside the test data, and thus also avoid deleting observations in the estimation data for trips that are in segments 1 and 10 (see the earlier point about data leakage).
Figure 1: Estimation and test data
Table 1: Estimation and test data sample sizes for case studies
@@
Distance (km) Estimation test Distance (km) Estimation test
@@
out of distribution for MA (below, i.e. $d_{min}-d_{a,MA}$ <1.33 0 0 223 <0.75 0 0 1,581
out-of-distribution for sub-models (below, i.e. $d_{a,MA}$ to $d_{a,sub}$) 1.33 to 1.95 0 836 261 0.75 to 1.11 0 6,446 1,642
within distribution for MA and sub-models 1.95 to 11.95 5,272 1,322 1.11 to 7.41 38,761 9,790
out-of-distribution for sub-models (above, i.e. $d_{b,sub}$ to $d_{b,MA}$) 11.95 to 19.52 0 946 153 7.41 to 11.22 0 6,458 1,646
out of distribution for MA (above, i.e. $d_{b,MA}tod_{max}$ 19.52 to 105.9 0 0 154 11.22 to 40.1 0 0 1,630
total 5,272 1,782 2,113 38,761 12,904 16,289
@@ 3.3 Model types
The following individual models are used in our analysis, combining structures from choice modelling, mathematical psychology and machine learning. Many other model types could have been considered [^1] A notable omission in our list is mixture models such as mixed logit or latent class - any advantages that such models have in estimation largely/completely disappear in prediction as the random heterogeneity needs to be averaged out. The only exception comes with the use of posterior distributions, which are not applicable in the context of out-of-sample prediction. - we limited ourselves to this set with a view to demonstrating the potential benefit of our proposed approach.
@@
Our NL models (cf. Daly and Zachary, 1978) use a full set (i.e. J1J-1) of alternative specific constants (ASC). For continuous attributes, we use a linear plus logarithmic specification, i.e. $V=...+\beta_{x}x+\beta_{log-x}log(x)$ to capture non-linearity in sensitivities. For the DECISIONS data, we use three nests, grouping car together with taxi, bus with train, and walking with cycling. For the LPMC data, we use two nests, grouping car together with bus, and walking together with cycling, but this structure ended up collapsing to a Multinomial Logit (MNL) model.
@@
This model comprises an input layer with input nodes, one or more hidden layers with hidden nodes, and an output layer with output nodes. In this model, signals propagate forward through the links connecting the nodes. The links have numeric weights $w$, which are learned from the data. At each link, the weights are multiplied by the input value from the previous nodes. At the node, the sum of the inputs is taken, and an activation function is applied. The outcome is propagated to the nodes in the next layer. We use tanh activation functions in all hidden layers, except the final one, where a linear layer is used. In the output layer, a Softmax function (i.e. a logit) is applied to produce choice probabilities for each alternative. To ensure consistency with the behavioural models and enhance predictive performance, availability conditions are explicitly incorporated into the network (thus, treated in the same way as in traditional choice models).
@@
The XGB model comprises a series of sequentially applied decision trees. A decision tree is a sequence of simple IF-THEN rules, optimised to classify data accurately. In the XGB model, each decision tree in the series ‘corrects’ the mispredictions of the models before it. This process is referred to as ‘boosting’. The term ‘gradient’ refers to the fact that each successive tree is fitted using the gradient of the loss function—effectively via gradient descent. Unlike standard Gradient Boosting Machines (GBM), XGBoost includes further optimisations, such as second-order gradient information and advanced regularisation, to improve speed and mitigate overfitting. To ensure consistency with the behavioural models and to better handle stronger regularisation settings, availability conditions are enforced via base_margin: we set margins to 0 for available alternatives and a large negative value for unavailable ones, ensuring that the model assigns zero probability to non-available alternatives. The model is implemented directly with the xgboost package.
@@
The NL and DFT models were coded and estimated using Apollo (cf. Hess and Palma, 2019), while the MLP and XGB models were implemented in Python. The MLP is implemented in PyTorch (Pedregosa et al., 2011) using a custom to handle the availability conditions, while the XGBoost model is implemented directly with the xgboost package (Chen and Guestrin, 2016). For both models, a hyperparameter grid search was conducted to optimise performance. We used a 5-fold Grouped cross-validation procedure to account for the panel structure of the data (see Table 2). The hyperparameter ranges for the XGB model were informed by the results of Hillel (2021).
@@
The number of hidden layers and nodes per layer influences the model’s capacity to learn complex patterns, while the batch size affects the stability and efficiency of the gradient updates. The learning rate controls how quickly the model updates its weights during training, while the L2 regularisation helps prevent overfitting by penalising large weight values. Since the best-performing hyperparameters were highly similar across both datasets, we opted to use a single configuration. The best-performing configuration found was two hidden layers with 30 nodes each, a learning rate of 0.001, a batch size of 250 and an L2 regularisation strength of 0. Because each training run of an MLP can yield slightly different outcomes, we trained the model 100 times using the best-performing configuration and averaged its predictions to account for this variability.
For the XGBoost model, the hyperparameter search included the maximum tree depth, the number of estimators, the minimum loss reduction, the minimum child weight, the subsample fraction, the fraction of features considered per split, regularisation strengths (L1 and L2), and the maximum delta step. The maximum tree depth controls the complexity of each tree, with deeper trees allowing more detailed partitions at the risk of overfitting. The number of estimators determines how many boosting rounds are performed, thereby influencing the overall capacity of the model. The minimum loss reduction acts as a pruning mechanism by preventing splits that do not provide sufficient improvement to the objective function, while the minimum child weight enforces a lower bound on the sum of instance weights required in a leaf, discouraging overly specific partitions. The subsample fraction regulates the proportion of training samples used for each boosting round, and the feature subsampling parameter (colsample by tree) increases model robustness by introducing randomness at the feature level. L1 and L2 regularisation terms penalise overly large weights and help stabilise the model under strong regularisation. Finally, the maximum delta step restricts the weight update size, improving convergence stability in cases of unbalanced data. Unlike the MLPs, the optimal hyperparameters for the XGBoost model were not identical across the datasets. Both the London and Decision data shared a fixed learning rate of 0.01, a minimum loss reduction of 0.6, a minimum child weight of 20, and a subsample fraction of 0.7. However, the London configuration used a feature subsampling rate of 0.75, a maximum tree depth of 6, 1,200 boosting rounds, and regularisation strengths of $L_{1}=20$ and $L_{2}=40$, with a maximum delta step of 6. In contrast, the Decision configuration relied on full feature sampling (1.0), a deeper maximum tree depth of 9, 800 boosting rounds, and lighter $L_{1}$ regularisation of 10 (with $L_{2}=40$), combined with a smaller maximum delta step of 4. As with the MLP model, we repeated the training 100 times using the optimal configuration and averaged their predictions to account for variability in the model training. This is common practice, as when training MLPs, a substantial share gets stuck in inferior solutions.
Table 2: Hyperparameter spaces considered for the MLP and XGBoost models.
MLP XGBoost
Hidden layers: $\{(10,5),(10,10),$ (20,20),(30,30)}\;(20,20),(30,30)\} Batch size: $\{250,500,1000\}$ Learning rate: $\{0.01,0.001,0.0001\}$ $L_{2}$ regularisation: $\{0,0.1,0.5,1\}$ Max. tree depth: $\{6,9\}$ Number of estimators: $\{800,1200\}$ Min. loss reduction: $\{0.2,0.6\}$ Min. child weight: $\{10,20\}$ Subsample fraction: $\{0.7,0.8\}$ Colsample by tree: $\{0.75,1.0\}$ $L_{1}$ regularisation: $\{10,20,40\}$ $L_{2}$ regularisation: $\{10,20,40\}$ Max. delta step: $\{4,6\}$
@@ 3.4 Model averaging: specification and estimation
Similar to the MLP models implemented for the training on the data sets (see Section 3.3), we implemented the MLP for the model averaging in PyTorch Paszke et al. (2019) too. This is needed because the loss function (see equation (2)) requires a custom implementation, which is not possible in e.g. SciKit-learn. As explanatory features, we included three distance-based variables: the raw trip distance, its logarithmic transformation, and its squared term. While in theory the linear distance alone should suffice, we found that adding the log- and squared transformations provided a small but consistent improvement in model performance. To tune the hyperparameters, we performed a grid search using a simple 75%–25% train–test split. Ideally, a dedicated validation set would have been used for this purpose—separate from the data employed in training the individual models, see Hastie (2009), who argues that not doing so might result in the meta-model assigning unfairly high weights to models with higher complexity. However, we deliberately refrained from further partitioning the data to avoid diminishing the already limited estimation sample and to ensure that the test set remained completely untouched for the final performance evaluation. The following parameter spaces are used: batch size (64, 128, 256), number of hidden layers and nodes (two or three hidden layers with each 10, 20, 40, or 60 nodes), learning rate (0.001, 0.005, 0.01) and L2 regularisation (0, 0.0001, 0.001). For the DECISIONS data set, the best-performing model has two hidden layers with 20 nodes each, a batch size of 128, a learning rate of 0.005, and an L2 regularisation strength of 0.001. For the LPMC data set, the best-performing model has two hidden layers with 60 nodes each, a batch size of 64, a learning rate of 0.001, and an L2 regularisation strength of 0.0001.
We repeated the training 100 times using the optimal configuration. As we aim for high prediction performance, we retained only the top 20 per cent of models [^2] This selection was based on the log-likelihood. For this, we used the full data, i.e. training and validation data, but excluded the test data., discarding those that became trapped in poor local maxima. This approach, thus, differs slightly from the one used for the training of the data-driven models, where we also repeated training 100 times but averaged predictions across all models to encourage diversity between models.
Figure 2 shows the structure for our model averaging approach. As shown in the figure, the model averaging MLP only takes the trip distance as input to predict the weights given to each sub-model. The model-averaging probability is computed as a weighted sum, combining the predictions for the trip by each sub-model and the weights assigned to each sub-model by the model-averaging MLP.